OpenStudy (anonymous):

A cylinder and a cone have the same diameter: 8 inches. The height of the cylinder is 3 inches. The height of the cone is 18 inches. Use π = 3.14. What is the relationship between the volume of this cylinder and this cone? Explain your answer by determining the volume of each and comparing them. Show all your work.

OpenStudy (anonymous):

I WILL MEDAL!!! i just need someone to show me steps on solving it because i almost have it solved but not quite

OpenStudy (anonymous):

Volume of cylinder:\[V=\pi r^{2}h\] Volume of cone: \[V=\pi r^{2} {h/3}\]

OpenStudy (anonymous):

this is what i have for my cylinder volume. cylinder v: 1/3*3.14*64*3=198.95

OpenStudy (anonymous):

idk if its right

OpenStudy (anonymous):

\[Radius=diameter/2\]

OpenStudy (anonymous):

Cylinder: \[V=(3.14)(4)^{2}(3)\]

OpenStudy (anonymous):

3.14*4*4*3=150.72. am i right?

OpenStudy (anonymous):

Cone: \[V=(3.14)(4)^{2}(18)(1/3)\]

OpenStudy (anonymous):

That's correct.

OpenStudy (anonymous):

so for the cone V its, 3.14*4*4*18*0.33=298.42. correct?

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

so what next?

OpenStudy (anonymous):

is that all?

OpenStudy (anonymous):

@chrismoon

OpenStudy (anonymous):

The relationship is that the volume of the cylinder is about half of the volume of the cone.

OpenStudy (anonymous):

is that all i have to do?

OpenStudy (anonymous):

@chrismoon thank you and i gave you a medal.

OpenStudy (anonymous):

also sorry for rushing i just got to get this done.

OpenStudy (anonymous):

I'd say that's all you have to do. The question doesn't ask you to do anything else..

OpenStudy (anonymous):

nope!